# Universal Vessel Segmentation for Multi-Modality Retinal Images

### v2

@@ III-C Topology-aware Segmentation Network
The downstream segmentation network is trained with a Perceptual Topological Loss $\mathcal{L}_{tc}$ and a Persistent Topological Loss $\mathcal{L}_{ts}$, introduced below accordingly.
@@ III-C1 Perceptual Topological Loss
We first propose to use a perceptual topological loss [59] where the topological information is approximated by the perceptual features encoded by an ImageNet [60] pretrained VGG-16 [27] network. As illustrated by Fig. 4, image embeddings $\mathcal{E}_{l}(\cdot)$ are computed from the predicted vessel probability map $I_{pred}$ and the ground truth $I_{GT}$, where $l$ denotes different layers in the VGG-16 network. In practice, we take the output of the $relu2\_2$, $relu3\_3$ and $relu4\_3$ layers. Then, the perceptual topological loss function $\mathcal{L}_{tc}$ is defined as:
$\mathcal{L}_{tc}=\sum_{l}^{N_{l}}\left|\mathcal{E}_{l}(I_{pred})-\mathcal{E}_{l}(I_{GT})\right|^{2}$ (6)
@@ III-C2 Persistent Topological Loss
The persistent homology of an image can be described by a persistent diagram, where each point on the diagram represent one persistent feature, whose coordinate is (1--Creation Value, 1--Destruction Value).
@@
$[\inf_{\eta:\mathcal{D}(I_{pred})\rightarrow\mathcal{D}(I_{GT})}\sum_{p\in\mathcal{D}(I_{pred})}s_{p}\left\|p-\eta(p))\right\|^{q}]^{\frac{1}{q}}$
@@ III-C4 Segmentation Network
We use a basic U-Net [28] structure with reduced model size for the segmentation network. In addition, as shown in Fig. 6, we adopt a ‘predict-refine’ strategy which first make a coarse prediction of the vessels and feed the initial prediction and the input image to the network again to refine it. This method has been shown useful in both topology-aware image segmentation [59] and multiple works in deep-learning-based retinal vessel segmentation [38, 39, 40]. In our method, the prediction and refinement networks share the same weight. Additionally, at both the initial prediction stage and the refinement stage, the loss functions are computed to optimize the network.
@@ IV Experiments
TABLE I: Comparison between our UVSM with SOTA finetuned methods. A green box indicates the best result and a blue box indicates the second-best result. Heidel refers to Heidelberg Spectralis.
Dataset Modality Camera #img Method Acc. Dice Sp Se Pr F1 MCC clDice $\beta_{0}\downarrow$ $\beta_{1}\downarrow$
$STARE$ $CF$ Topcon 20 W-Net 97.61 81.52 98.80 81.59 82.36 81.52 80.51 85.56 75.6 13.1
@@
$ChaseDB1$ $CF$ Zeiss 28 W-Net 97.08 78.87 98.75 76.35 82.32 78.87 77.58 78.46 54.09 9.0
@@
$HRF$ $CF$ Canon 35 W-Net 96.60 64.39 98.48 61.70 68.04 64.39 62.88 60.09 260.4 59.8
WS-DMF 94.79 58.02 96.02 71.78 49.75 58.03 56.84 56.54 945.1 51.7
U-Net 96.58 64.22 98.42 61.82 67.40 64.22 62.64 59.63 297.7 53.1
CS2-Net 96.68 65.17 98.51 62.37 68.82 65.18 63.67 60.82 334.7 51.1
FA-Net 96.67 65.80 98.40 64.36 67.87 65.80 64.23 61.97 257.7 47.5
Our UVSM 96.60 68.50 97.78 74.70 63.87 68.50 67.15 68.26 142.8 55.5
$IOSTAR$ $MC$ Heidel 30 W-Net 97.09 80.66 98.70 78.54 83.50 80.66 79.30 87.46 43.2 9.4
WS-DMF 95.89 71.49 98.29 67.53 77.54 71.49 69.88 76.92 140.9 22.0
GT-DLA-dsHff 96.91 79.63 98.52 78.27 81.67 79.63 78.16 86.99 43.4 7.8
U-Net 97.01 80.50 98.52 79.72 81.80 80.50 79.04 87.08 57.4 8.7
CS2-Net 97.05 80.59 98.59 79.27 82.49 80.59 79.17 87.06 67.2 10.3
FA-Net 97.13 81.38 98.52 81.30 82.23 81.38 80.06 88.08 38.3 10.9
Our UVSM 96.70 78.38 98.36 77.53 79.72 78.38 76.73 85.08 69.1 6.3
$JRCFA$ $FA$ Heidel Optos 40 W-Net 97.83 83.49 99.03 81.73 86.11 83.49 82.57 84.68 162.0 13.7
WS-DMF 97.20 78.20 98.69 76.90 81.84 78.20 77.35 78.60 209.2 14.8
U-Net 97.89 83.95 99.14 81.65 87.02 83.95 83.04 84.86 160.5 13.9
CS2-Net 97.98 84.88 99.00 84.43 85.79 84.88 83.93 86.07 157.6 11.2
FA-Net 97.94 84.47 99.03 83.51 85.94 84.47 83.51 85.72 126.3 10.1
Our UVSM 96.47 75.00 97.91 78.40 72.89 74.89 73.43 76.37 37.7 15.2
$JRCFAF$ $FAF$ Heidel Optos 40 W-Net 97.93 77.78 99.21 74.01 82.59 77.77 76.98 81.36 48.0 8.1
WS-DMF 97.48 72.19 99.05 67.88 78.94 72.19 71.53 75.63 109.1 10.6
GT-DLA-dsHff 97.24 70.71 98.80 67.71 74.57 70.71 69.50 74.61 62.8 7.5
U-Net 97.89 77.66 99.13 74.94 81.29 77.66 76.81 81.03 68.6 7.2
CS2-Net 97.93 78.11 99.14 75.34 81.64 78.12 77.24 81.46 63.4 6.9
FA-Net 97.95 78.59 99.09 76.69 81.17 78.69 77.70 82.55 42.4 7.3
Our UVSM 97.30 75.49 97.96 85.35 68.19 75.49 74.78 81.89 21.6 4.5
$JRCIR$ IR NIR Heidel 40 W-Net 97.07 79.79 98.85 75.25 83.35 78.80 77.52 79.19 55.3 7.0
WS-DMF 96.32 74.49 98.10 74.37 75.87 74.49 72.88 75.67 107.9 9.8
GT-DLA-dsHff 96.30 73.68 98.30 71.36 76.64 73.68 71.88 74.46 88.5 5.9
U-Net 97.14 79.92 98.66 78.74 81.84 79.92 78.60 80.93 79.4 5.6
CS2-Net 97.22 80.28 98.77 78.30 83.01 80.28 79.01 80.79 86.9 5.6
FA-Net 97.24 80.48 98.75 78.79 82.84 80.48 79.19 81.30 49.6 5.0
Our UVSM 96.10 74.66 97.84 75.31 72.85 73.66 71.81 76.30 28.8 5.5
Figure 8: Qualitative comparison between our UVSM and finetuned SOTA models on multi-modality retinal images (better zoom in to view). Red, blue orange and green box include the original image, the ground truth label, finetuned FA-Net segmentation and the segmentation from our UVSM. From left to right, up to down: Topcon CF, Canon CF, Zeiss CF, Heidelberg MC, Optos FA, Heidelberg FA, Optos FAF, Heidelberg FAF, Heidelberg IR. Bottom right shows the qualitative results for Optos MC and Canon FA (no ground truth available to compare with and to train the baseline finetuned models).
@@ IV-A1 Image Translation Training
We prepared a multi-modality database with over 150k de-identified images from the database of the Jacobs Retina Center at UCSD. However, in our previous numerous attempts we found it very difficult and time-consuming to tune the training of the image translation network (regardless of the methods) for a large dataset. Therefore, in this work, we use a 4k images (2k CF, 500 FA, FAF, MC, IR/NIR, respectively) dataset from the larger 150k database. Each image has a resolution of 768 $\times$ 768. All the CF images are from the 30-degree (field-of-view) Topcon camera and all the IR/NIR images are from the Heidelberg camera. For FA, FAF and MC, 250 images are from Heidelberg camera. For Optos camera, the original images are ultra-wide-filed 135-degree images, which are out of the scope of this work (will be dicussed in section V-A). Therefore, we crop random regions from the center area of the retina to get 30-40-degree conventional field-of-view images and apply a commonly-used circular mask (Fig. 7(a)). In the dataset, there are 250 cropped and masked Optos image for FA, MC, FAF modalities, respectively.
@@ IV-A2 Segmentation Training
We train the segmentation network using only the 20 training images from the public DRIVE [10] CF dataset. In practice, we discard image #14 since the vessel pattern is unusual. In addition, for all the images we resize them to 768$\times$768 for training. We also resized the ground truth annotation and manually corrected the error caused by interpolation. In addition, according to our examination, the original DRIVE ground truth has multiple topological errors such as small redundant holes in the vessel foreground. In our revised ground truth, we correct these errors and allow learning more accurate vessel topological features.
@@ IV-A3 Evaluation
Seven diverse datasets are used for evaluation. For the CF modality, we use the public STARE [78], ChaseDB1 [79] and HRF [16] datasets. For MC modality, we use the public IOSTAR [80] dataset. For FA, FAF and IR/NIR, we prepared three new datasets: JRCFA, JRCFAF and JRCIR. Each dataset has 40 768$\times$768 images, which is more than most of the commonly-used public retinal vessel segmentation (CF) datasets. The JRCFA and JRCFAF datasets both consist 20 images from Heidelberg and 20 from Optos, respectively. The JRCIR dataset consists 40 Heidelberg (the only camera for this modality) images. All the vessel annotations are either labeled or carefully examined and refined by the retina experts in the UCSD Jacobs Retina Center (JRC). More information on the 7 datasets can be found in Table I. In addition to the seven vessel segmentation dataset, we also evaluate our model qualitatively on the Optos MC images using images from our 150K database and on the Canon FA images on the CF-FA registration dataset [81]. We provide preliminary insights in this paper (Fig. 1, Fig. 8) yet full evaluation requires the vessel labels (unavailable) and is left as future work.

### v3

@@ III-C Topology-aware Segmentation Network
The downstream segmentation network is learned with a novel combination of Perceptual Topological Loss $\mathcal{L}_{tc}$ and a Persistent Topological Loss $\mathcal{L}_{ts}$, introduced below accordingly.
@@ III-C1 Perceptual Topological Loss
We first propose to use a perceptual topological loss [62] where the topological information is approximated by the perceptual features encoded by an ImageNet [63] pretrained VGG-16 [30] network. As illustrated by Fig. 4, image embeddings $\mathcal{E}_{l}(\cdot)$ are computed from the predicted vessel probability map $I_{pred}$ and the ground truth $I_{GT}$, where $l$ denotes different layers in the VGG-16 network. In practice, we take the output of the relu2_2relu2\_2, relu3_3relu3\_3 and relu4_3relu4\_3 layers. Then, the perceptual topological loss function $\mathcal{L}_{tc}$ is defined as:
$\mathcal{L}_{tc}=\sum_{l}^{N_{l}}\left|\mathcal{E}_{l}(I_{pred})-\mathcal{E}_{l}(I_{GT})\right|^{2}_{2}$ (6)
@@ III-C2 Persistent Topological Loss
The persistent homology of an image can be described by a persistent diagram, where each point on the diagram represents one persistent feature, whose coordinate is (1-Creation Value, 1-Destruction Value).
@@
$[\inf_{\eta:\mathcal{D}(I_{pred})\rightarrow\mathcal{D}(I_{GT})}\sum_{p\in\mathcal{D}(I_{pred})}s_{p}\left\|p-\eta(p))\right\|^{q}_{2}]^{\frac{1}{q}}$
@@ III-C4 Segmentation Network
We use a basic U-Net [31] structure with reduced model size for the segmentation network. In addition, as shown in Fig. 6, we adopt a ‘predict-refine’ strategy which first makes a coarse prediction of the vessels and feeds the initial prediction and the input image to the network again to refine it. This method has been shown useful in both topology-aware image segmentation [62] and multiple works in deep-learning-based retinal vessel segmentation [41, 42, 43]. In our method, the prediction and refinement networks share the same weight. Additionally, at both the initial prediction stage and the refinement stage, the loss functions are computed to optimize the network.
@@ IV Experiments
TABLE I: Details of evaluation datasets
Dataset modality camera No. images
STARE CF Topcon 20
HRF CF Canon 35
ChasedDB1 CF Zeiss 28
IOSTAR MC Heidelberg 30
JRCFA FA Heidelberg & Optos 40
JRCFAF FAF Heidelberg & Optos 40
JRCIR IR Heidelberg 40
OCTA500-3M OCTA Optovue 40
OCTA500-6M OCTA Optovue 100
PRIME-FP20 MC(UWF) Optos 15
Figure 8: Qualitative comparison between our UVSM and finetuned SOTA models on multi-modality retinal images (better zoom in to view). Red, blue orange and green box include the original image, the ground truth label, finetuned FA-Net segmentation and the segmentation from our UVSM. From left to right, up to down: Topcon CF, Canon CF, Zeiss CF, Heidelberg MC, Optos FA, Heidelberg FA, Optos FAF, Heidelberg FAF, Heidelberg IR. Optovue OCTA 3mm, Optovue OCTA 6mm, Bottom right shows the qualitative results for Optos MC and Canon FA (no ground truth available to compare with and to train the baseline finetuned models).
TABLE II: Comparison between our UVSM with SOTA finetuned (unsupervised domain adaptation and supervised) methods on Non - Color Fundus modality (excluding OCTA). The green boxes indicate the best result and the blue boxes indicate the second-best result.
[…]
 IOSTARIOSTAR (MC)  DomainDomain  AdaptationAdaptation Pse. Lab. 96.27 74.53 98.42 71.10 79.35 74.53 72.92 80.55 75.3 9.1
FRR-TSNet 96.25 73.37 98.72 67.39 81.93 73.37 72.11 79.11 73.1 10.1
MLAL 95.33 72.85 96.57 80.85 67.14 72.85 70.99 77.61 180.1 9.8
MLAL-PDSF 95.73 74.58 96.98 81.12 69.37 74.58 72.65 80.04 138.3 6.3
SupervisedSupervised W-Net 97.09 80.66 98.70 78.54 83.50 80.66 79.30 87.46 43.2 9.4
WS-DMF 95.89 71.49 98.29 67.53 77.54 71.49 69.88 76.92 140.9 22.0
GT-DLA-dsHff 96.91 79.63 98.52 78.27 81.67 79.63 78.16 86.99 43.4 7.8
U-Net 97.01 80.50 98.52 79.72 81.80 80.50 79.04 87.08 57.4 8.7
CS2-Net 97.05 80.59 98.59 79.27 82.49 80.59 79.17 87.06 67.2 10.3
FA-Net 97.13 81.38 98.52 81.30 82.23 81.38 80.06 88.08 38.3 10.9
Our UVSM 96.70 78.38 98.36 77.53 79.72 78.38 76.73 85.08 69.1 6.3
 JRCFAJRCFA (FA)  DomainDomain  AdaptationAdaptation Pse. Lab. 88.10 6.05 94.15 5.93 7.61 6.05 0.36 3.55 251.7 34.6
FRR-TSNet 92.71 0.60 99.56 0.33 6.01 6.04 -0.29 0.38 53.0 30.4
MLAL 85.25 5.42 91.07 6.25 5.45 5.42 -2.07 3.02 450.1 32.2
MLAL-PDSF 86.47 4.86 92.49 5.25 5.43 4.86 -1.88 2.73 411.6 23.5
SupervisedSupervised W-Net 97.83 83.49 99.03 81.73 86.11 83.49 82.57 84.68 162.0 13.7
WS-DMF 97.20 78.20 98.69 76.90 81.84 78.20 77.35 78.60 209.2 14.8
U-Net 97.89 83.95 99.14 81.65 87.02 83.95 83.04 84.86 160.5 13.9
CS2-Net 97.98 84.88 99.00 84.43 85.79 84.88 83.93 86.07 157.6 11.2
FA-Net 97.94 84.47 99.03 83.51 85.94 84.47 83.51 85.72 126.3 10.1
Our UVSM 96.47 75.00 97.91 78.40 72.89 74.89 73.43 76.37 37.7 15.2
 JRCFAFJRCFAF (FAF)  DomainDomain  AdaptationAdaptation Pse. Lab. 94.26 63.03 94.76 84.39 52.93 63.03 63.34 66.55 438.9 30.3
FRR-TSNet 97.52 74.34 98.73 73.91 76.60 74.34 73.53 76.86 136.5 7.6
MLAL 95.21 64.64 95.69 85.59 53.70 64.64 64.99 64.45 482.7 10.9
MLAL-PDSF 95.66 66.66 96.21 85.09 56.58 66.66 66.75 67.13 425.4 19.8
SupervisedSupervised W-Net 97.93 77.78 99.21 74.01 82.59 77.77 76.98 81.36 48.0 8.1
WS-DMF 97.48 72.19 99.05 67.88 78.94 72.19 71.53 75.63 109.1 10.6
GT-DLA-dsHff 97.24 70.71 98.80 67.71 74.57 70.71 69.50 74.61 62.8 7.5
U-Net 97.89 77.66 99.13 74.94 81.29 77.66 76.81 81.03 68.6 7.2
CS2-Net 97.93 78.11 99.14 75.34 81.64 78.12 77.24 81.46 63.4 6.9
FA-Net 97.95 78.59 99.09 76.69 81.17 78.69 77.70 82.55 42.4 7.3
Our UVSM 97.30 75.49 97.96 85.35 68.19 75.49 74.78 81.89 21.6 4.5
 JRCIRJRCIR (IR)  DomainDomain  AdaptationAdaptation Pse. Lab. 95.91 71.89 97.84 71.85 73.17 71.89 70.05 71.15 86.9 8.3
FRR-TSNet 96.06 71.95 98.21 69.38 76.13 71.95 70.30 71.02 112.1 15.3
MLAL 94.84 69.21 96.26 76.85 64.93 69.21 67.44 68.55 239.4 14.1
MLAL-PDSF 95.31 70.07 97.00 74.25 68.91 70.07 68.43 69.12 205.6 11.2
SupervisedSupervised W-Net 97.07 79.79 98.85 75.25 83.35 78.80 77.52 79.19 55.3 7.0
WS-DMF 96.32 74.49 98.10 74.37 75.87 74.49 72.88 75.67 107.9 9.8
GT-DLA-dsHff 96.30 73.68 98.30 71.36 76.64 73.68 71.88 74.46 88.5 5.9
U-Net 97.14 79.92 98.66 78.74 81.84 79.92 78.60 80.93 79.4 5.6
CS2-Net 97.22 80.28 98.77 78.30 83.01 80.28 79.01 80.79 86.9 5.6
FA-Net 97.24 80.48 98.75 78.79 82.84 80.48 79.19 81.30 49.6 5.0
Our UVSM 96.10 74.66 97.84 75.31 72.85 73.66 71.81 76.30 28.8 5.5
TABLE III: Comparison between our UVSM with SOTA finetuned (unsupervised domain adaptation and supervised) methods on Color Fundus modality. The green boxes indicate the best result and the blue boxes indicate the second-best result.
[…]
 STARESTARE  DomainDomain  AdaptationAdaptation Pse. Lab. 97.17 76.06 98.99 70.97 84.07 76.06 75.36 80.71 71.7 19.0
FRR-TSNet 97.15 74.41 99.10 68.72 85.88 74.41 74.44 79.31 64.1 19.4
MLAL 95.85 69.96 97.17 76.01 68.06 69.96 68.95 73.30 122.1 15.3
MLAL-PDSF 96.39 73.12 97.76 76.73 71.89 73.12 71.88 76.09 129.2 15.6
SupervisedSupervised W-Net 97.61 81.52 98.80 81.59 82.36 81.52 80.51 85.56 75.6 13.1
@@
 HRFHRF  DomainDomain  AdaptationAdaptation Pse. Lab. 97.00 67.36 98.83 62.73 74.10 67.36 66.38 60.19 69.6 71.5
FRR-TSNet 97.07 66.27 99.12 58.55 78.75 66.27 66.03 59.22 75.1 74.1
MLAL 95.65 61.96 96.94 71.07 55.87 61.96 60.52 60.77 103.8 59.0
MLAL-PDSF 95.93 63.37 97.26 70.95 57.89 63.37 61.81 61.71 95.0 60.3
SupervisedSupervised W-Net 96.60 64.39 98.48 61.70 68.04 64.39 62.88 60.09 260.4 59.8
WS-DMF 94.79 58.02 96.02 71.78 49.75 58.03 56.84 56.54 945.1 51.7
U-Net 96.58 64.22 98.42 61.82 67.40 64.22 62.64 59.63 297.7 53.1
CS2-Net 96.68 65.17 98.51 62.37 68.82 65.18 63.67 60.82 334.7 51.1
FA-Net 96.67 65.80 98.40 64.36 67.87 65.80 64.23 61.97 257.7 47.5
Our UVSM 96.60 68.50 97.78 74.70 63.87 68.50 67.15 68.26 142.8 55.5
 $ChaseDB1$  DomainDomain  AdaptationAdaptation Pse. Lab. 96.07 71.24 98.25 68.35 75.18 71.25 69.44 70.05 68.6 6.8
FRR-TSNet 95.94 68.93 98.53 62.93 76.63 68.93 67.26 67.17 83.5 18.3
MLAL 94.99 68.96 96.30 78.01 62.07 68.96 66.89 67.44 203.1 23.2
MLAL-PDSF 95.59 71.17 97.06 76.61 66.73 71.17 69.08 69.93 146.3 18.1
SupervisedSupervised W-Net 97.08 78.87 98.75 76.35 82.32 78.87 77.58 78.46 54.09 9.0
@@
TABLE IV: Performance on Ultra-Wide-Field retinal images, pure zero-shot performance on OCTA, and comparison with SOTA supervised methods.
[…]
OCTA5003MOCTA500-3M (OCTA) UNet++ 95.98 88.64 - -
FARGO 98.12 91.68 - -
FRNet-base 98.95 91.29 99.50 90.56
FRNet 99.00 91.70 99.51 91.19
Our UVSM 96.29 74.68 97.36 81.97
OCTA5006MOCTA500-6M (OCTA) UNet++ 95.73 85.76 - -
FARGO 98.12 89.15 - -
FRNet-base 98.11 88.82 99.05 88.01
FRNet 98.14 89.03 99.01 88.77
Our UVSM 95.21 71.68 98.20 66.53
[…]
PRIMEFP20PRIME-FP20 (UWF-MC) LKM-UNet 58.11 41.00 57.55 51.55
Swin-Umamba 64.99 48.43 64.36 61.93
EM-Net 62.30 45.31 61.84 55.40
Serp-Mamba 69.01 52.93 68.41 65.75
Our UVSM 65.40 49.00 65.69 74.32
TABLE V: Comparison between UVSM and general-purpose foundational segmentation models, reporting metric: Dice score.
CF MC FA FAF IR OCTA
SAM2 15.45 18.19 16.12 12.05 17.08 15.80
Grounded SAM2 17.57 18.11 15.77 12.04 17.10 16.34
MedSAM2 14.74 16.99 9.68 6.90 6.66 14.58
Our UVSM 76.21 78.38 75.00 75.49 74.66 72.54
@@ IV-A1 Image Translation Training
We prepared a multi-modality database with over 150k de-identified images from the database of the Jacobs Retina Center at UCSD. However, in our previous numerous attempts we found it very difficult and time-consuming to tune the training of the image translation network (regardless of the methods) for a large dataset. Therefore, in this work, we use a 4k images (2k CF, 500 FA, FAF, MC, IR/NIR, respectively) dataset from the larger 150k database. Each image has a resolution of 768 $\times$ 768. All the CF images are from the 30-degree (field-of-view) Topcon camera and all the IR/NIR images are from the Heidelberg camera. For FA, FAF and MC, 250 images are from Heidelberg camera. For Optos camera, we crop random regions from the original 135 degrees to get 30-40-degree conventional field-of-view images (so that the image size is the same as the other modalities) and apply a commonly-used circular mask (Fig. 7(a)). In the dataset, there are 250 cropped and masked Optos image for FA, MC, FAF modalities, respectively.
@@ IV-A2 Segmentation Training
We train the segmentation network using only the 20 training images from the public DRIVE [10] CF dataset. In practice, we discard image #14 due to the unusual vessel pattern. In addition, for all the images we resize them to $\times768$ for training. We also resized the ground truth annotation and manually corrected the error caused by interpolation. In addition, according to our examination, the original DRIVE ground truth has multiple topological errors such as small redundant holes in the vessel foreground. In our revised ground truth, we correct these errors and allow more efficient topological learning.
@@ IV-A3 Evaluation
Ten diverse datasets are used for evaluation. Their details can be found in Table I. For the CF modality, we use the public STARE [83], ChaseDB1 [84] and HRF [19] datasets. For MC modality, we use the public IOSTAR [85] dataset. For FA, FAF and IR/NIR, we prepared three new datasets: JRCFA, JRCFAF and JRCIR. Each dataset has 40 $\times768$ images, which is more than most of the commonly-used public retinal vessel segmentation (CF) datasets. The JRCFA and JRCFAF datasets both consist 20 images from Heidelberg and 20 from Optos, respectively. The JRCIR dataset consists 40 Heidelberg (the only camera for this modality) images. All the vessel annotations are either labeled or carefully examined and refined by the retina experts in the UCSD Jacobs Retina Center (JRC). Furthermore, we use an OCTA500-3M and an OCTA600-6M datasets [86] to evaluate the performance of the UVSM on OCTA in two different field-of-views (3mm and 6mm). For the special UWF image (technically UWF is not a modality but only refer to the field-of-view), we use a PRIME-FP20 (MC) dataset [87]. In addition to the ten vessel segmentation dataset, we also evaluate our model qualitatively on the Optos MC images using images from our 150K database and on the Canon FA images on the CF-FA registration dataset [88]. We provide preliminary insights in this paper (Fig. 1, Fig. 8) yet full evaluation requires the vessel labels (unavailable) and is left as future work.